# PREPRINT VERSION

### v3

@@ V-B1 Denoising DINOv2 Features
We employ the denoised version of DINOv2 as described in [23] and highlight its effectiveness as a feature extractor through the results shown in Table III. We present a comparison between the original [22] and denoised [23] DINOv2 deep feature embeddings in Fig. 7. Additionally, we compare the performance of DINOv2 trained with registers [60] to the denoised version trained with registers [60, 23], though this approach did not yield any improvement.
@@
Figure 7: A comparison is made between Point Label Aware Superpixels (PLAS) [4] features, raw DINOv2 features [22], DINOv2 features trained with registers [60], denoised DINOv2 features [23], and denoised DINOv2 features trained with registers [23, 60] for images in the UCSD Mosaics dataset. In the case of the transformer-based methods, features for each 14x14 pixel patch in the original image are upsampled with bilinear interpolation. Features are visualized in RGB by reducing them with Principal Component Analysis (PCA), where the first three components represent the R, G and B channels. Pixels visualized in similar colors indicate similarity in the deep embedding space. The CNN features used by Point Label Aware Superpixels (PLAS) [4] fail to cluster pixels into distinct segments which align with the expected ground truth segments. The denoised model significantly reduces artifacts from position embeddings, leading to smoother, cleaner features and better clustering performance. In the coral imagery context, the models trained with registers do not seem to improve the feature space.
@@
Figure 8: Features extracted from ViTs are subjected to noisy grid-like artifacts caused by the positional encoding in the architecture. Patch tokens are extracted from the DINOv2 [22] pre-trained ViT, spatially upsampled with bilinear interpolation and then visualized with random projection into the RGB colorspace. Sub-figure a) shows the query images, which are deliberately constant black/white images to demonstrate the artifacts, b) shows the features with noise artifacts from the positional encoding, c) shows that the features are free from noise when the positional encoding is not used, and d) shows that the Denoising ViT [23] significantly reduces the noise artifacts. Positional encoding is needed during training to inform the spatial relationships between the image patches, enabling the model to effectively learn structure and context in images.
@@ V-B2 Weighting the Probability Maps ($\lambda$)
We assess how the $\lambda$ weight, which adjusts the significance of the cosine similarity map (Eq. 2) relative to the distance map (Eq. 3), affects performance. As illustrated in Fig. 9, a $\lambda$ value of 2.2 yields the highest pixel accuracy. However, our method is relatively insensitive to the precise value of $\lambda$. The analysis indicates that variations in pixel accuracy are less than 1% across a broad range of $\lambda$ values tested ($1.5\leq\lambda\leq 3$).
@@ V-B3 Exclusion Distance ($\sigma$)
The proposed human-in-the-loop labeling regime accounts for the proximity of labeled pixels by using a Gaussian-smoothed distance mask that measures the distance from all pixels to those that have already been labeled. The Gaussian smoothing introduces the $\sigma$ hyperparameter, which dictates how near new labeled points can be to existing point labels (Eq. 4). Fig. 9 presents an ablation study on the effect of the $\sigma$ hyperparameter on the performance of point label propagation, and we find that our method remains effective across various values of $\sigma$, ranging from 1 to 100.
@@ V-D Effect of Point Label Placement Style
VI Conclusion
This study has shown that the general foundation model, DINOv2, can be effectively utilized for point label propagation in underwater imagery without requiring fine-tuning. By leveraging denoised DINOv2 features and a straightforward KNN algorithm, we generate augmented ground truth masks. When combined with our human-in-the-loop labeling approach, we achieve significant improvements in mIoU: 19.7%, 18.3%, and 8.2% for 5, 10, and 25 point labels, respectively, when compared to previous state-of-the-art methods. Even when using DINOv2 features and KNN with grid-spaced point labels, we outperform earlier methods on the UCSD dataset with improvements of 5.8%, 10.0%, and 4.6% mIoU for 5, 10, and 25 point labels, respectively.
These performance gains are evident in the semantic segmentation task as well. Training a DeepLabv3+ model on augmented ground truth masks created with DINOv2, KNN, and our human-in-the-loop labeling approach yields improvements of 8.8% in pixel accuracy and 13.5% in mIoU with just 5 point labels. If the human-in-the-loop regime is not employed, the semantic segmentation still benefits, showing an 6.5% increase in pixel accuracy and a 6.2% increase in mIoU.
Our comprehensive studies on the impact of the number of point labels and their placement reveal that grid labels consistently enhance point label propagation accuracy compared to random labels. For more than 100 points per image, improvements in point label propagation are minimal with either the Point Aware Superpixels [4] method or our DINOv2 and KNN approach. However, for very few points per image, domain expert input in selecting which pixels to label proves advantageous. This work highlights the effectiveness of general foundation models for complex, domain-specific tasks and significantly boosts performance and annotation efficiency in scenarios with extremely sparse labels.
@@ VI Conclusion
The large number of coral species, the complexity of coral imagery, and the location-specific nature of coral visual features mean that real-world deployment of multi-species image analysis systems must efficiently gather data in the target environment, label a few points, and then train and deploy a model. With an estimated 800 species of hard corals alone [61], collecting and annotating sufficient data for a global model is impractical. Instead, we believe that efforts should focus on rapidly training and deploying models tailored to specific locations. The method proposed in this work contributes to this goal by significantly reducing the amount of required annotations while achieving accurate point label propagation.
@@ Acknowledgments
This work was done in collaboration between QUT and CSIRO Data61. S.R., F.M., N.S., and T.F. acknowledge continued support from the Queensland University of Technology (QUT) through the Centre for Robotics. T.F. acknowledges funding from Intel Labs via grant RV3.290.Fischer and an ARC Laureate Fellowship FL210100156 to Prof. Michael Milford. Computational resources and services used in this work were provided by the eResearch Office, Queensland University of Technology, Brisbane, Australia.

### v4

@@ V-B1 Denoising DINOv2 Features
We employ the denoised version of DINOv2 as described in [23] and highlight its effectiveness as a feature extractor through the results shown in Table III. We present a comparison between the original [22] and denoised [23] DINOv2 deep feature embeddings in Fig. 10. Additionally, we compare the performance of DINOv2 trained with registers [27] to the denoised version trained with registers [27, 23], though this approach did not yield any improvement. The denoising with registers approach slightly outperforms the denoising version for mean pixel accuracy only in the 5 and 10 point label cases, and for all other metrics and label quantities there is no improvement from the addition of registers.
@@
Figure 10: A comparison is made between Point Label Aware Superpixels (PLAS) [4] features, raw DINOv2 features [22], DINOv2 features trained with registers [27], denoised DINOv2 features [23], and denoised DINOv2 features trained with registers [23, 27] for images in the UCSD Mosaics dataset. In the case of the transformer-based methods, features for each 14x14 pixel patch in the original image are upsampled with bilinear interpolation. Features are visualized in RGB colors by reducing them with Principal Component Analysis (PCA), where the first three components represent the R, G and B channels respectively. Pixels visualized in similar colors indicate similarity in the deep embedding space. The CNN features used by Point Label Aware Superpixels (PLAS) [4] fail to cluster pixels into distinct segments which align with the expected ground truth segments. The denoised model significantly reduces artifacts from position embeddings, leading to smoother, cleaner features and better clustering performance. In the coral imagery context, the models trained with registers do not seem to improve the feature space. Refer to Fig. 6 for the color legend.
@@
Figure 11: Features extracted from ViTs are subjected to noisy grid-like artifacts caused by the positional encoding in the architecture. Patch tokens are extracted from the DINOv2 [22] pre-trained ViT, upsampled and visualized into RGB colors using the first three components after reduction with PCA, following the same procedure as in Fig. 10. Sub-figure a) shows the query images, which are deliberately constant black/white images to demonstrate the artifacts, as well as an example coral image from UCSD Mosaics [24, 2], b) shows the features with noise artifacts from the positional encoding, c) shows that the features are free from noise when the positional encoding is not used, and d) shows that the Denoising ViT [23] significantly reduces the noise artifacts. Positional encoding is needed during training to inform the spatial relationships between the image patches, enabling the model to effectively learn structure and context in images.
@@ V-B2 Weighting the Probability Maps ($\lambda$)
We assess how the $\lambda$ weight, which adjusts the significance of the cosine similarity map (2) relative to the distance map (3), affects performance. As illustrated in Fig. 12, a $\lambda$ value of 2.2 yields the highest pixel accuracy. However, our method is relatively insensitive to the precise value of $\lambda$. The analysis indicates that variations in pixel accuracy are less than 1% across a broad range of $\lambda$ values tested ($1.5\leq\lambda\leq 3$).
@@ V-B3 Exclusion Distance ($\sigma$)
The proposed human-in-the-loop labeling regime accounts for the proximity of labeled pixels by using a Gaussian-smoothed distance mask that measures the distance from all pixels to those that have already been labeled. The Gaussian smoothing introduces the $\sigma$ hyperparameter, which dictates how near new labeled points can be to existing point labels (4). Fig. 12 presents an ablation study on the effect of the $\sigma$ hyperparameter on the performance of point label propagation, and we find that our method remains effective across various values of $\sigma$, ranging from 1 to 100.
@@ V-B5 Effect of Initial Human-Labeled Points
V-B6 Effect of Stage 2 Segmentation Architecture and Backbone
In this section, we investigate the performance of different semantic segmentation architectures and backbones. For consistency, we train the models all on our propagated masks, generated using our HIL approach. Table IV compares the performance of the different architectures UNet [63] (comprised of 5 blocks), LinkNet [64] (with ResNet50 [61] backbone), and Deeplabv3+ [60] on the test dataset. We also compare various backbone architectures for the Deeplabv3+ architecture (ResNet50 [61], Inceptionv3 [62] and MobileNetv2 [69]).
TABLE IV: Performance of Stage Two: Comparison of Semantic Segmentation Architectures and Backbones (Refer to Section IV-C for Metric Definitions).
Architecture Backbone PA mPA mIoU Inference Time (fps) UNet [63] n/a 80.40 49.30 40.65 16.74 LinkNet [64] ResNet50 [61] 68.85 44.84 36.88 12.99 Deeplabv3+ [60] MobileNetv2 [69] 78.43 45.65 38.60 13.33 Deeplabv3+ [60] Inceptionv3 [62] 78.74 51.85 41.28 6.70 Deeplabv3+ [60] ResNet50 [61] 74.53 41.47 32.96 13.84
The results in Table IV demonstrate that the Deeplabv3+ with Inceptionv3 backbone outperforms UNet, LinkNet (with ResNet50), and DeepLabv3+ (with ResNet50 and MobileNetv2 backbones) for coral image segmentation. This is likely due to Inceptionv3’s multi-scale convolutional filters within its inception modules, which enable it to capture both local textures and broader contextual information effectively. ResNet50 primarily relies on deep residual connections and sequential convolutions, and MobileNetv2 prioritizes efficiency with depthwise separable convolutions, meaning that Inceptionv3 maintains comparably richer, more diverse feature maps. It is likely that UNet is the next highest performing architecture due to its strong preservation of fine-grained spatial details. UNet’s symmetric encoder-decoder structure, with skip connections at multiple scales, allows it to retain and refine high-resolution features, which is crucial for separating semantically similar coral and marine invertebrate classes. While LinkNet also incorporates skip connections, it has a lightweight decoder which is less capable of restoring intricate boundaries compared to UNet’s more expansive upsampling path.
For consistency with prior approaches, the Deeplabv3+ architecture with ResNet50 backbone is used to evaluate the impact of different point label propagation approaches .
Despite the Deeplabv3+ with Inceptionv3 backbone achieving higher performance on the semantic segmentation task, the Deeplabv3+ architecture with ResNet50 backbone is used to evaluate the impact of different point label propagation approaches (Table II), to maintain consistency with prior approaches [3, 4].
@@ V-D Effect of Point Label Placement Style
VI Discussion
VI-A Impact
The core impact of this work is in reducing the cost associated with expert annotation of coral reef imagery. Labeling images for semantic segmentation is expensive, as domain experts must manually annotate each point, with costs accumulating per labeled point. By significantly decreasing the number of required point labels while maintaining or improving segmentation accuracy, our approach leads to direct cost savings. For example, if traditional methods require 300 point labels per image and our method achieves comparable performance with only 5 point labels, this represents a 98.3% reduction in labeling effort and therefore cost. Typically, datasets should be in the thousands of images to enable sufficient generalization to a range of visual conditions – this translates into significant financial savings, making model fine-tuning and adaptation for large-scale coral reef monitoring more cost-effective and feasible.
This paper presents a comprehensive set of experiments analyzing how the quantity of point labels influences segmentation performance. Given that the definition of satisfactory segmentation is subjective and varies based on user needs, we evaluate multiple labeling schemes across different levels of human input. Our results illustrate the trade-offs between annotation effort and segmentation accuracy, showing that while additional point labels generally improve performance, highly accurate point label propagation with fewer point labels can still be obtained by using our proposed approach (Fig. 8). The number of point labels should be selected based on the desired use case, the dataset characteristics, labeling time and budget, and the availability of expert annotators. Rather than prescribing a fixed number of necessary labels, we provide insights into how different annotation strategies impact outcomes, allowing users to make informed decisions based on their specific constraints and application requirements.
VI-B Limitations
Our approach is limited by the two stages in the architecture i.e. one stage of processing to generate the augmented ground truth masks for the training data, followed by another stage for training the segmentation model on the augmented masks. It would be more efficient to employ a single stage approach which takes as input the underwater image and sparse point labels, and outputs dense segmentation masks at inference time. A single stage approach was trialled in which deep features for labelled pixels were stored in a memory bank and used as templates for clustering pixels in new imagery. However, the vision transformer architecture encodes global information in each embedding, meaning that pixels of the same class in different images are far apart in the deep embedding space. A single-stage approach which overcomes this limitation is left for future work.
The human-in-the-loop labeling regime in this work was simulated using the ground truth masks with the UCSD Mosaics dataset [24, 2]. The regime starts by the ‘domain expert’ labelling up to ten pixels centrally in coral instances, and then iterative prediction of the optimal locations for labelling the next points up to a pre-defined maximum quantity of point labels. This was simulated by obtaining the ground truth label at the identified pixel location as though it was labelled by an expert ‘in-the-loop’. The approach would benefit from user testing with multiple domain experts to determine how inter-observer variability impacts the performance of the approach.
VI-C Future Work
@@
VII Conclusion
Marine ecosystem monitoring requires efficient analysis of vast coral reef imagery, but expert annotation for semantic segmentation is prohibitively expensive and time-consuming, creating a critical bottleneck for large-scale ecological studies. This work has demonstrated for the first time how general foundation models can be effectively utilized for point label propagation in domain-specific underwater imagery without requiring fine-tuning, combined with a novel strategic human-in-the-loop labeling framework that selects annotation locations based on feature uncertainty and spatial diversity.
By leveraging denoised DINOv2 features with a straightforward KNN algorithm, we generate augmented ground truth masks from extremely sparse labels (5-25 points per image) through a two-stage pipeline that first propagates point labels and then trains semantic segmentation models on the resulting augmented masks. When combined with our human-in-the-loop labeling approach, we achieve significant improvements in mIoU for the label propagation task: 19.7%, 18.3%, and 8.2% for 5, 10, and 25 point labels, respectively, when compared to previous state-of-the-art methods. These performance gains are evident in the semantic segmentation task as well: Training a DeepLabv3+ model on augmented ground truth masks created with DINOv2, KNN, and our human-in-the-loop labeling approach yields improvements of 8.8% in pixel accuracy and 13.5% in mIoU with just 5 point labels.
The large number of coral species (there are an estimated 800 species of hard corals alone [72]), visual complexity of coral imagery, and geographic diversity of species mean that real-world deployment of underwater computer vision must involve data collection in the target environment, efficient annotation, and then model training or fine-tuning. The method proposed in this work contributes to the goal of rapid training and deployment of location-specific models by significantly reducing the amount of required annotations while achieving accurate point label propagation, with broader implications for cost-effective computer vision in specialized domains where expert annotation is scarce and expensive.
@@ Acknowledgments
This work was done in collaboration between QUT and CSIRO Data61. S.R., F.M., N.S., and T.F. acknowledge continued support from the Queensland University of Technology (QUT) through the Centre for Robotics. T.F. acknowledges funding from an Australian Research Council (ARC) Discovery Early Career Researcher Award (DECRA) Fellowship DE240100149. S.R. and T.F. acknowledge support from the Reef Restoration and Adaptation Program (RRAP) which is funded by a partnership between the Australian Government’s Reef Trust and the Great Barrier Reef Foundation. Computational resources and services used in this work were provided by the eResearch Office, Queensland University of Technology, Brisbane, Australia.